The Grand Hyatt Manila located in Taguig City is currently the highest building in the Philippines with a height of approximately 318 m. Suppose a 65-kg runner runs up the fire exit stairs of the building from the ground floor up to its rooftop in 15 minutes. Determine the runners power output in kilowatts and in horsepower. 


1
Expert's answer
2021-02-24T12:49:24-0500
"P=\dfrac{W}{t}=\dfrac{mgh}{t},""P=\dfrac{65\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot318\ m}{15\cdot60\ s}=0.225\ kW,""P=\dfrac{65\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot318\ m}{15\cdot60\ s}=225\ W\cdot\dfrac{1\ hp}{746\ W}=0.3\ hp."